1.) Two numbers have a sum of 3 and a product of -40. Find the numbers.




2.) Two numbers have a sum of -6 and a product of -16. Find the numbers.

Answer by algebrahouse.com(1659)

1.) Two numbers have a sum of 3 and a product of -40. Find the numbers.
= 8,-5
8(-5) = -40
8 + (-5) = 3

2.) Two numbers have a sum of -6 and a product of -16. Find the numbers.
= -8,2
-8(2) = -16
-8 + 2 = -6
www.algebrahouse.com